# Quantile Activation: departing from single point estimation for better generalization across distortions

### v1

@@ Abstract
A classifier is, in its essence, a function which takes an input and returns the class of the input and implicitly assumes an underlying distribution. We argue in this article that one has to move away from this basic tenet to obtain generalization across distributions. Specifically, the class of the sample should depend on the points from its “context distribution” for better generalization across distributions.
How does one achieve this? – The key idea is to “adapt” the outputs of each neuron of the network to its context distribution. We propose quantile activation,QAct, which, in simple terms, outputs the relative quantile of the sample in its context distribution, instead of the actual values in traditional networks.
The scope of this article is to validate the proposed activation across several experimental settings, and compare it with conventional techniques. For this, we use the datasets developed to test robustness against distortions – CIFAR10C, CIFAR100C, MNISTC, TinyImagenetC, and show that we achieve a significantly higher generalization across distortions than the conventional classifiers, across different architectures. Although this paper is only a proof of concept, we surprisingly find that this approach outperforms DINOv2(small) at large distortions, even though DINOv2 is trained with a far bigger network on a considerably larger dataset.
@@ 1 Introduction
Deep learning approaches have significantly influenced image classification tasks on the machine learning pipelines over the past decade. They can easily beat human performance on such tasks by non-trivial margins by using innovative ideas such as Batch Normalization [19] and other normalization techniques [3, 31], novel rectifiers such as ReLU/PReLU [26, 31, 3] and by using large datasets and large models.
However, these classification systems do not generalize across distributions [2, 34], which leads to instability when used in practice. [22] shows that deep networks with ReLU activation degrades in performance under distortions. [4] observes that there exists a feature collapse which inhibits the networks to be reliable.
Fundamental Problem of Classification:
We trace the source of the problem to the fact that – Existing classification pipelines necessitates single point prediction, i.e., they should allow classification of a single sample given in isolation. We argue that, for good generalization, one should move away from this basic tenet and instead allow the class to be dependent on the “context distribution” of the sample. That is, when performing classification, one needs to know both the sample and the context of the sample for classification.
While this is novel in the context of classification systems, it is widely prevalent in the specific domain of Natural Language Processing (NLP) – The meaning of a word is dependent on the context of the word. However, to our knowledge, this has not been considered for general classification systems. Even when using dominant NLP architectures such as Transformers for vision [8], the technique has been to split the image into patches and then obtain the embedding for an individual image.
@@
Obtaining a classification framework for incorporating context distribution:
We suspect that the main reason why the context distribution is not incorporated into the classification system is – The naive approach of considering a lot of samples in the pipeline to classify a single sample is computationally expensive. We solve this problem by considering the context distribution of each neuron specifically. We introduce Quantile Activation (QAct), which outputs a probability depending upon the context of all outputs. This, however, gives rise to new challenges in training, which we address in section 2.
Figure 1 illustrates the differences of the proposed framework with the existing framework. As severity increases (w.r.t Gaussian Noise), we observe that ReLU activation loses the class structure. This behaviour can be attributed to the fact that, as the input distribution changes, the activations either increase/decrease, and due to the multiplicative effect of numerous layers, this leads to very different features. On the other hand, the proposed QAct framework does not suffer from this, since if all the pre-activations [^1] We use the following convention – “Pre-activations” denote the inputs to the activation functions and “Activations” denote the outputs of the activation function. change in a systematic way, the quantiles adjust automatically to ensure that the inputs for the next layer does not change by much. This is reflected in the fact that class structure is preserved with QAct.
Remark:
Quantile activation is different from existing quantile neural network based approaches, such as regression [30], binary quantile classification [36], Anomaly Detection [24, 33]. Our approach is achieving best in-class performance by incorporating context distribution in the classification paradigm. Our approach is also markedly different from Machine unlearning which is based on selective forgetting of certain data points or retraining from scratch [32].
@@
    Our approach contributes to achieving better generalization across distributions and is more robust to distortions, across architectures. We evaluate our method using different architectures and datasets, and compare with the current state-of-the-art – DINOv2. We show that QAct proposed here is more robust to distortions than DINOv2, even if we have considerably less number of parameters (22M for DINOv2 vs 11M for Resnet18). Additionally, DINOv2 is trained on 20 odd datasets, before being applied on CIFAR10C; in contrast, our framework is trained on CIFAR10, and produces more robust outcome (see figures 3,5).
@@
    The proposed QAct is consistent with all the existing techniques used in DINOv2, and hence can be easily incorporated into any ML framework.
  5. 5.
    We also adapt QAct to design a classifier which returns better calibrated probabilities.
Related Works on Domain Generalization (DG):
The problem of domain generalizations tries to answer the question – Can we use a classifier trained on one domain across several other related domains? The earliest known approach for this is Transfer Learning [28, 37], where a classifier from a single domain is applied to a different domain with/without fine-tuning. Several approaches have been proposed to achieve DG, such as extracting domain-invariant features over single/multiple source domains [11, 1, 9, 29, 16], Meta Learning [17, 9], Invariant Risk Minimization [2]. Self supervised learning is another proposed approach which tries to extract features on large scale datasets in an unsupervised manner, the most recent among them being DINOv2 [27]. Very large foundation models, such as GPT-4V, are also known to perform better with respect to distribution shifts [12]. Nevertheless, to the best of our knowledge, none of these models incorporates context distributions for classification.
@@ 2 Quantile Activation
An alternate way to model a neuron is to consider it as predicting a latent variable ${\mathbf{y}}$, where ${\mathbf{y}}=1$ if the feature is present and ${\mathbf{y}}=0$ if the feature is absent. Mathematically, we have the following model:
${\mathbf{z}}=w^{t}{\mathbf{x}}+b+\epsilon\quad and\quad{\mathbf{y}}=I[{\mathbf{z}}\geq 0]$ (1)
This is very similar to the standard latent variable model for logistic regression, with the main exception being, the outputs ${\mathbf{y}}$ are not known for each neuron beforehand. If ${\mathbf{y}}$ is known, it is rather easy to obtain the probabilities – $P({\mathbf{z}}\geq 0)$. Can we still predict the probabilities, even when ${\mathbf{y}}$ itself is a latent variable?
The authors in [5] propose the following algorithm to estimate the probabilities:
@@
    Let $\{{\bm{x}}_{i}\}$ denote the set of input samples from the input distribution ${\mathbf{x}}$ and $\{z_{i}\}$ denote their corresponding latent outputs, which would be from the distribution ${\mathbf{z}}$
@@
    Assign ${\mathbf{y}}=1$ whenever ${\mathbf{z}}>(1-\tau)^{th}$ quantile of ${\mathbf{z}}$, and 00 otherwise. For a specific sample, we have $y_{i}=1$ if $z_{i}>(1-\tau)^{th}$ quantile of $\{z_{i}\}$
@@
Observe that in step 2., the labelling is done without resorting to actual ground-truth labels. This allows us to obtain the probabilities on the fly for any set of parameters, only by considering the quantiles of ${\mathbf{z}}$.
@@
Let ${\mathbf{z}}$ denote the pre-activation of the neuron, and let $\{{\bm{z}}_{i}\}$ denote the samples from this distribution. Let $F_{{\mathbf{z}}}$ denote the cumulative distribution function (CDF), and let $f_{{\mathbf{z}}}$ denote the density of the distribution. Accordingly, we have that $F_{{\mathbf{z}}}^{-1}(\tau)$ denotes the $\tau^{th}$ quantile of ${\mathbf{z}}$. Using step (2) of the algorithm above, we define,
@@
However, to use QAct in a neural network, we need to compute the gradient which is required for back-propagation. Let $\tau_{{\bm{z}}}$ denote the quantile at which $F_{{\mathbf{z}}}^{-1}(\tau_{{\bm{z}}})={\bm{z}}$. Then we have that $\textsc{QAct}{}({\bm{z}})=\tau_{z}$ since $F_{{\mathbf{z}}}^{-1}(\tau)$ is an increasing function. So, we have that $\textsc{QAct}{}(F_{{\mathbf{z}}}^{-1}(\tau))=\tau$. In other words, we have that $\textsc{QAct}{}({\bm{z}})$ is $F_{{\mathbf{z}}}({\bm{z}})$, which is nothing but the CDF of ${\mathbf{z}}$. Hence, we have,
$\frac{\partial\textsc{QAct}{}({\bm{z}})}{\partial z}=f_{{\mathbf{z}}}({\bm{z}})$ (4)
where $f_{{\mathbf{z}}}({\bm{z}})$ denotes the density of the distribution.
@@
  Obtain a kernel density estimate, using points from $S$, at each of the points in $z_{i}$ – (say) $\hat{f_{{\mathbf{z}}}}(z_{i})$
@@
With the above formulation, observe that since QAct is identical to CDF, it follows that, $\textsc{QAct}{}({\mathbf{z}})$ is always a uniform distribution between 00 and $1$, irrespective of the distribution ${\mathbf{z}}$. When training numerous neurons in a layer, this could cause all the neurons to learn the same behaviour. Specifically, if, half the time, a particular abstract feature is more prevalent than others, QAct (as presented above) would not be able to learn this feature. To correct this, we enforce that positive values and negative values have equal weight. Given the input distribution ${\mathbf{z}}$, We perform the following transformation before applying QAct. Let
${\mathbf{z}}^{+}=\begin{cases}{\mathbf{z}}&\text{ if }{\mathbf{z}}\geq 0\\ 0&\text{ otherwise}\end{cases}\hskip 80.00012pt{\mathbf{z}}^{-}=\begin{cases}{\mathbf{z}}&\text{ if }{\mathbf{z}}<0\\ 0&\text{ otherwise}\end{cases}$ (7)
@@
${\mathbf{z}}^{\ddagger}=\begin{cases}{\mathbf{z}}^{+}&\text{ with probability }0.5\\ {\mathbf{z}}^{-}&\text{ with probability }0.5\\ \end{cases}$ (8)
From definition of ${\mathbf{z}}^{\ddagger}$, we get that the median of ${\mathbf{z}}^{\ddagger}$ is 00. This grounds the input distribution to have the same positive and negative weight.
@@
Note that the gradient for the back propagation is given by the density of ${\mathbf{z}}^{\ddagger}$ (weighted distribution). We use the Kernel Density Estimation (KDE), to estimate the density. We, (i) First sample $N$ points with weights $w_{+},w_{-}$, and (ii) then estimate the density at all the input points $[{\bm{z}}_{i}]$. This is point-wise multiplied with the backward gradient to get the gradient for the input. In this article we use $N=1000$, which we observe gets reasonable estimates.
@@
Computational Complexity (for a single neuron) is majorly decided by 2 functions – (i) Computing the quantiles has the complexity for a vector $[{\bm{z}}_{i}]$ of size $n$ can be performed in ${\mathcal{O}}(n\log(n))$. Since this is log-linear in $n$, it does not increase the complexity drastically compared to other operations in a deep neural network. (ii) Computational complexity of the KDE estimates is ${\mathcal{O}}(Sn_{\tau})$ where $S$ is the size of sample (weighted sample from $[{\bm{z}}_{i}]$) and $n_{\tau}$ is the number of quantiles, giving a total of ${\mathcal{O}}(n+Sn_{\tau})$. In practice, we consider $S=1000$ and $n_{\tau}=100$ which works well, and hence does not increase with the batch size. This too scales linearly with batch size $n$, and hence does not drastically increase the complexity.
@@
To understand the idea behind quantile activation, consider a simple toy example in figure 2. For ease of visualization, assume that the input features (blue) are in 2 dimensions, and also assume that the line of the linear projection is given by the red line in figure 2(a). Now, assume that the blue input features are rotated, leading to a different distribution (indicated here by orange). Since activations are essentially (unnormalized) signed distances from the line, we plot the histograms corresponding to the two distributions in figure 2(b). As expected, these distributions are different. However, after performing the quantile activation in equation 3, we have that both are uniform distribution. This is illustrated in figures 2(c) and 2(d). This behaviour has a normalizing effect across different distributions, and hence has better distribution generalization than other activations.

### v2

# Quantile Activation: Correcting a failure mode of ML models

@@ Abstract
An established failure mode for machine learning models occurs when the same features are equally likely to belong to class 00 and class $1$. In such cases, existing ML models cannot correctly classify the sample. However, a solvable case emerges when the probabilities of class 00 and $1$ vary with the context distribution. To the best of our knowledge, standard neural network architectures like MLPs or CNNs are not equipped to handle this.
In this article, we propose a simple activation function, quantile activation (QAct), that addresses this problem without significantly increasing computational costs. The core idea is to “adapt” the outputs of each neuron to its context distribution. The proposed quantile activation, QAct, produces the relative quantile of the sample in its context distribution, rather than the actual values, as in traditional networks.
A practical example where the same sample can have different labels arises in cases of inherent distribution shift. We validate the proposed activation function under such shifts, using datasets designed to test robustness against distortions—CIFAR10C, CIFAR100C, MNISTC, TinyImagenetC. Our results demonstrate significantly better generalization across distortions compared to conventional classifiers, across various architectures. Although this paper presents a proof of concept, we find that this approach unexpectedly outperforms DINOv2 (small) under large distortions, despite DINOv2 being trained with a much larger network and dataset.
@@ 1 Introduction
Thanks to deep learning approaches, machine learning has been adopted across wide variety of domains. However, there is a significant failure mode within the standard framework of machine learning – A specific sample can potentially be labelled 00 or $1$ depending on the context. Hence, standard ML models cannot handle these conflicts effectively.
A simple toy example to illustrate the failure mode of ML systems:
We use a simple toy example to describe this failure mode. Consider the distribution generated as follows (figure 1(a)):
$\mu_{1}$ $\sim\mathcal{U}(S^{1})$    random sample from uniform distribution on the circle (1)
$\mu_{2}$ $=R(30^{\circ})\mu_{1}$ $\quad\mu_{2}\text{ is obtained by rotating }\mu_{1}\text{ by }30^{\circ}$
Class 0 $\sim\mathcal{N}(\mu_{1},0.1{\bm{I}})$ Class 0 generated using normal with mean $\quad\text{Class }0\text{ generated using normal with mean }\mu_{1}\text{ and stdev }0.1$
Class 1 $\sim\mathcal{N}(\mu_{2},0.1{\bm{I}})$ Class 1 generated using normal with mean $\quad\text{Class }1\text{ generated using normal with mean }\mu_{2}\text{ and stdev }0.1$
Why is this classification problem hard? Note that any point $x$ from the support of the above distribution is equally likely to be class 00 or class $1$. So, one cannot construct any fixed function depending only on the input features $x$. Thus, most of existing ML frameworks, which insist on learning a fixed function, fail. Nevertheless, this is a valid distribution where classification with accuracy $1$ is theoretically possible when one can reconstruct the latent $\mu_{1}$, $\mu_{2}$.
Specifically, the current neural network architectures fail as well. Consider training a simple MLP on this dataset using gradient descent. That is, each batch (of size $B$) of samples is generated as - Sample $\mu_{1},\mu_{2}$, then sample $B/2$ points each from $\mathcal{N}(\mu_{1},0.1{\bm{I}})$ and $\mathcal{N}(\mu_{2},0.1{\bm{I}})$ respectively. Since, we would have that a specific sample is equally likely to belong class 00 or $1$, one would learn (probability) $p=0.5$ for all the samples. This is verified in figure 1(b) where we use ReLU activation.
Remark:
We remark that the failure mode described above is one of the possible failure modes, but there can exist other failure modes as well. In this article, we are interested in rectifying the above failure mode alone.
Fixing the failure mode and defining the “context”:
The only way to fix the above failure mode is to infer the $\mu_{1},\mu_{2}$ from the batch of samples. We define context as the batch of samples according to which each sample is processed. To our knowledge current neural network architectures such as MLP/CNN do not consider this. While transformers considers this to some extent via the self-attention module, it is still sample specific and is very expensive computationally to obtain self-attention for the entire batch. Moreover, vision transformers do not consider attention across different samples in the batch.
Quantile Activation can identify the context distribution from each batch:
To rectify this shortcoming, we propose quantile activation, QAct, in this article. The key idea is that – at each neuron, the final activation value depends on the activations of the entire batch of samples. Specifically, we use the relative quantile of the pre-activation  [^1] We use the following convention – “Pre-activations” denote the inputs to the activation functions and “Activations” denote the outputs of the activation function. with respect to other pre-activations in the batch. (Details in section 2). Figure 1(b) shows that this simple change can allow the neural network to learn in spite of the contradictory labels.
Figure 1: A simple toy example to illustrate where ML systems fail. (a) The distribution is a mixture of Gaussian distributions whose centers $(\mu_{1},\mu_{2})$ are separated by $30^{\circ}$. The centres themselves can lie anywhere on the unit circle. (please refer to the text for exact description). (b) Histogram of accuracy over $1000$ different combinations of $\mu_{1},\mu_{2}$ for both ReLU activation and after incorporating QAct. Clearly, ReLU activation alone cannot perform better than random guess. Incorporating QAct on the other hand can easily infer the latent $\mu_{1},\mu_{2}$.
@@
Distribution Shift as a failure mode:
The inability of current network architectures to handle distribution shift is a manifestation of the failure mode discussed above. For instance consider the above distribution (figure 1(a)), where, if the pair of Gaussian distributions are rotated by $30^{\circ}$ (a.k.a distribution shift), the class 00 and class $1$ overlap. We validate the proposed quantile activation on the standard distribution shift dataset - CIFAR10C. Figure 2 illustrates the results obtained using ReLU activations and QAct. As severity increases (w.r.t Gaussian Noise), we observe that ReLU activation loses the class structure. On the other hand, the proposed QAct framework does not suffer from this and the class structure is preserved with QAct.
Remark 1:
Although QAct deals with activation of neurons, it is fundamentally different compared to the usual activation functions such as ReLU/Gelu etc. QAct is designed for classification tasks where context plays an important role and hence considering “pure” classification tasks (i.e where the underlying distribution and ground-truth are unknown but fixed) is out of scope in this article.
Remark 2:
Quantile activation is different from existing quantile neural network based approaches, such as regression (Prashanth et al., 2022), binary quantile classification (Tambwekar et al., 2022), Anomaly Detection (Li & van Leeuwen, 2023; Seo et al., 2022). Our approach is achieving best in-class performance by incorporating context distribution in the classification paradigm. Our approach is also markedly different from Machine unlearning which is based on selective forgetting of certain data points or retraining from scratch (Seetha et al., 2024).
@@
    Our approach contributes to achieving better generalization across distributions and is more robust to distortions, across architectures. We evaluate our method using different architectures and datasets, and compare with the current state-of-the-art – DINOv2. We show that QAct proposed here is more robust to distortions than DINOv2, even if we have considerably less number of parameters (22M for DINOv2 vs 11M for Resnet18). Additionally, DINOv2 is trained on 20 odd datasets, before being applied on CIFAR10C; in contrast, our framework is trained on CIFAR10, and produces more robust outcome (see figures 4,6).
@@
    We also adapt QAct to design a classifier which returns better calibrated probabilities. We show that, unlike the relevant, most recent baselines (RELU, DINOv2 (small)), QAct achieves constant calibration error across different severity of distortions.
Related Works:
This work aims to address the failure mode described earlier. To the best of our knowledge, no existing literature specifically addresses this issue. However, we believe this issue to be widely persistent in several practical domains. We use distribution shift to validate the proposed approach, which is part of a bigger problem of domain generalization.
Related Works on Domain Generalization (DG): The problem of domain generalization tries to answer the question – Can we use a classifier trained on one domain across several other related domains? The earliest known approach for this is Transfer Learning (Pan & Yang, 2010; Zhuang et al., 2021), where a classifier from a single domain is applied to a different domain with/without fine-tuning. Several approaches have been proposed to achieve DG, such as extracting domain-invariant features over single/multiple source domains (Ghifary et al., 2015; Akuzawa et al., 2019; Dou et al., 2019; Piratla et al., 2020; Hu et al., 2019), Meta Learning (Huang et al., 2020; Dou et al., 2019), Invariant Risk Minimization (Arjovsky et al., 2019). Self-supervised learning is another proposed approach which tries to extract features on large scale datasets in an unsupervised manner, the most recent among them being DINOv2 (Oquab et al., 2023) which is the current state-of-the-art333as per https://paperswithcode.com/sota/domain-generalization-on-imagenet-c accessed on 26 September 2024. Very large foundation models, such as GPT-4V, are also known to perform better with respect to distribution shifts (Han et al., 2023). Nevertheless, to the best of our knowledge, none of these models incorporates context distributions for classification.
@@ 2 Quantile Activation
An alternate way to model a neuron is to consider it as predicting a latent variable y, where ${\textnormal{y}}=1$ if the feature is present and ${\textnormal{y}}=0$ if the feature is absent. Mathematically, we have the following model:
${\textnormal{z}}=w^{t}{\mathbf{x}}+b+\epsilon\quad and\quad{\textnormal{y}}=I[{\textnormal{z}}\geq 0]$ (2)
This is very similar to the standard latent variable model for logistic regression, with the main exception being, the outputs y are not known for each neuron beforehand. If y is known, it is rather easy to obtain the probabilities – $P({\textnormal{z}}\geq 0)$. Can we still predict the probabilities, even when y itself is a latent variable?
The authors in (Challa et al., 2023) propose the following algorithm to estimate the probabilities:
@@
    Let $\{{\bm{x}}_{i}\}$ denote the set of input samples from the input distribution ${\mathbf{x}}$ and $\{z_{i}\}$ denote their corresponding latent outputs, which would be from the distribution z
@@
    Assign ${\textnormal{y}}=1$ whenever ${\textnormal{z}}>(1-\tau)^{th}$ quantile of z, and 00 otherwise. For a specific sample, we have $y_{i}=1$ if $z_{i}>(1-\tau)^{th}$ quantile of $\{z_{i}\}$
@@
Observe that in step 2., the labelling is done without resorting to actual ground-truth labels. This allows us to obtain the probabilities on the fly for any set of parameters, only by considering the quantiles of z.
@@
Let z denote the pre-activation of the neuron, and let $\{z_{i}\}$ denote the samples from this distribution. Let $F_{{\textnormal{z}}}$ denote the cumulative distribution function (CDF), and let $f_{{\textnormal{z}}}$ denote the density of the distribution. Accordingly, we have that $F_{{\textnormal{z}}}^{-1}(\tau)$ denotes the $\tau^{th}$ quantile of z. Using step (2) of the algorithm above, we define,
@@
However, to use QAct in a neural network, we need to compute the gradient which is required for back-propagation. Let $\tau_{z}$ denote the quantile at which $F_{{\textnormal{z}}}^{-1}(\tau_{z})=z$. Then we have that $\textsc{QAct}{}(z)=\tau_{z}$ since $F_{{\textnormal{z}}}^{-1}(\tau)$ is an increasing function. So, we have that $\textsc{QAct}{}(F_{{\textnormal{z}}}^{-1}(\tau))=\tau$. In other words, we have that $\textsc{QAct}{}(z)$ is $F_{{\textnormal{z}}}(z)$, which is nothing but the CDF of z. Hence, we have,
$\frac{\partial\textsc{QAct}{}(z)}{\partial z}=f_{{\textnormal{z}}}(z)$ (5)
where $f_{{\textnormal{z}}}(z)$ denotes the density of the distribution.
@@
  Obtain a kernel density estimate, using points from $S$, at each of the points in $z_{i}$ – (say) $\hat{f_{{\textnormal{z}}}}(z_{i})$
@@
[…]
${\textnormal{z}}^{+}=\begin{cases}{\textnormal{z}}&\text{ if }{\textnormal{z}}\geq 0\\ 0&\text{ otherwise}\end{cases}\hskip 80.00012pt{\textnormal{z}}^{-}=\begin{cases}{\textnormal{z}}&\text{ if }{\textnormal{z}}<0\\ 0&\text{ otherwise}\end{cases}$ (8)
@@
${\textnormal{z}}^{\ddagger}=\begin{cases}{\textnormal{z}}^{+}&\text{ with probability }0.5\\ {\textnormal{z}}^{-}&\text{ with probability }0.5\\ \end{cases}$ (9)
From definition of ${\textnormal{z}}^{\ddagger}$, we get that the median of ${\textnormal{z}}^{\ddagger}$ is 00. This grounds the input distribution to have the same positive and negative weight.
@@
Note that the gradient for the back propagation is given by the density of ${\textnormal{z}}^{\ddagger}$ (weighted distribution). We use the Kernel Density Estimation (KDE), to estimate the density. We, (i) First sample $S$ points with weights $w_{+},w_{-}$, and (ii) then estimate the density at all the input points $[z_{i}]$. This is point-wise multiplied with the backward gradient to get the gradient for the input. In this article we use $S=1000$, which we observe gets reasonable estimates.
@@
Computational Complexity (for a single neuron) is majorly decided by 2 functions – (i) Computing the quantiles has the complexity for a vector $[z_{i}]$ of size $n$ can be performed in ${\mathcal{O}}(n\log(n))$. Since this is log-linear in $n$, it does not increase the complexity drastically compared to other operations in a deep neural network. (ii) Computational complexity of the KDE estimates is ${\mathcal{O}}(Sn_{\tau})$ where $S$ is the size of sample (weighted sample from $[z_{i}]$) and $n_{\tau}$ is the number of quantiles, giving a total of ${\mathcal{O}}(n\log(n)+Sn_{\tau})$. In practice, we consider $S=1000$ and $n_{\tau}=100$ which works well, and hence does not increase with the batch size.
@@
To understand the idea behind quantile activation, consider a simple toy example in figure 3. For ease of visualization, assume that the input features (blue) are in 2 dimensions, and also assume that the line of the linear projection is given by the red line in figure 3(a). Now, assume that the blue input features are rotated, leading to a different distribution (indicated here by orange). Since activations are essentially (unnormalized) signed distances from the line, we plot the histograms corresponding to the two distributions in figure 3(b). As expected, these distributions are different. However, after performing the quantile activation in equation 4, we have that both are uniform distribution. This is illustrated in figures 3(c) and 3(d). This behaviour has a normalizing effect across different distributions, and hence has better distribution generalization than other activations.